### v1

# On the sample complexity of estimation in logistic regression

Abstract
The logistic regression model is one of the most popular data generation model in noisy binary classification problems. In this work, we study the sample complexity of estimating the parameters of the logistic regression model up to a given $\ell_{2}$ error, in terms of the dimension and the inverse temperature, with standard normal covariates. The inverse temperature controls the signal-to-noise ratio of the data generation process. While both generalization bounds and asymptotic performance of the maximum-likelihood estimator for logistic regression are well-studied, the non-asymptotic sample complexity that shows the dependence on error and the inverse temperature for parameter estimation is absent from previous analyses. We show that the sample complexity curve has two change-points (or critical points) in terms of the inverse temperature, clearly separating the low, moderate, and high temperature regimes.
1 Introduction
This paper studies the sample complexity of estimating the parameter vector in the logistic regression model under a normal design, with particular attention paid to the dependence on the dimension $d$, the inverse temperature $\beta\geq 0$, and the target error $\epsilon\in\lparen 0,1\rparen$ ). We show how the form of the sample complexity changes depending on the particular relationship between the inverse temperature and the target error.
[…]
We assume the observed data $(\mathbf{x}_{1},\mathbf{y}_{1}),\dotsc,(\mathbf{x}_{n},\mathbf{y}_{n})$ are independent copies of $(\mathbf{x},\mathbf{y})$, with distribution determined by the unknown parameter vector $\theta^{\star}\in S^{d-1}$ (and inverse temperature $\beta$). For a given $\epsilon\in\lparen 0,1\rparen$ ), the goal is to find an estimate ${\hat{\theta}}={\hat{\theta}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})\in S^{d-1}$ based on these data (and possibly also $\beta$) such that ${\hat{\theta}}$ satisfies
$\lVert{\hat{\theta}}-\theta^{\star}\rVert\leq\epsilon$
(either in expectation or with high probability over the realization of the data).
The sample complexity $n^{*}(d,\beta,\epsilon)$ is the smallest sample size $n$ such that the above task is achievable by some estimator. We show that, up to logarithmic factors in $1/\epsilon$ in the expressions below, the sample complexity satisfies
$n^{*}(d,\beta,\epsilon)\asymp\begin{cases}\displaystyle\frac{d}{\beta^{2}\epsilon^{2}}&\text{if $\beta\lesssim 1$ (high temperatures)};\\ \displaystyle\frac{d}{\beta\epsilon^{2}}&\text{if $1\lesssim\beta\lesssim 1/\epsilon$ (moderate temperatures)};\\ \displaystyle\frac{d}{\epsilon}&\text{if $\beta\gtrsim 1/\epsilon$ (low temperatures)}.\end{cases}$
1.1 Motivation
Our original motivation for studying this problem comes from the application to noisy one-bit (compressive) sensing, in which only a single bit is retained per linear measurement of a signal (Boufounos and Baraniuk, 2008). In that context, Plan and Vershynin (2012) proposed a robust linear estimator that is well-behaved under a variety of observation models, including the logistic regression model that we consider. This estimator (which in the present context is essentially the same as the “Average” algorithm of Servedio (1999)) was shown by Plan et al. (2017) to have sample complexity
$O\left\lparen\frac{d}{\min\{\beta^{2},1\}\epsilon^{2}}\right\rparen,$
improving on an earlier analysis of Plan and Vershynin (2012) that had a worse dependence on $\epsilon$; note that here we do not assume that $\theta^{\star}$ is sparse. The optimality of this estimator in the high-temperature regime ($\beta\lesssim 1$) is readily established as a standard application of Fano’s inequality (see, e.g., Chen et al., 2016, Appendix C.1). However, it was unclear whether this sample complexity is optimal in other regimes. In particular, in the zero-temperature regime (i.e., $\beta=\infty$), the $\mathbf{y}_{i}$ are deterministically given by the sign of $\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}$, so the problem becomes equivalent to that of PAC learning homogeneous linear classifiers under spherically symmetric distributions on $\mathbf{x}$. For that problem, the sample complexity is $\Theta(d/\epsilon)$, as established by Long (1995, 2003); note that the dependence on $\epsilon$ is considerably reduced. [^1] Servedio (1999) studied the “Average” algorithm in this context, allowing for the possibility that each observed label is independently flipped with some fixed probability $\eta\in[0,1/2\rparen$ ), and obtains a sample complexity upper bound of $O(d/((1-2\eta)\epsilon)^{2})$. Jacques et al. (2013) also gives a statement of the lower bound in the context of one-bit compressive sensing.
Lower bounds on the sample complexity in our setting do not directly follow from standard lower bounds from statistical learning theory for homogeneous linear classifiers (e.g., Devroye and Lugosi, 1995), as the distributions exhibiting the lower bounds generally do not conform to our statistical model. In particular, the support of $\mathbf{x}$ in these lower bounds is typically taken to be a shattered finite set of points, and the conditional distribution of $\mathbf{y}$ given $\mathbf{x}$ may not be of the form in logistic regression. The exceptions, as mentioned above, are those based on Fano’s inequality for the high temperature regime, and the lower bound in the zero temperature regime of Long (1995).
Our upper and lower bounds resolve the dependence of the sample complexity on the inverse temperature (up to logarithmic factors in $1/\epsilon$), particularly in the regime where $1\lesssim\beta<\infty$.
1.2 Techniques
Our lower bounds for the high and moderate temperature regimes (Theorem 1) are proved using Fano’s inequality; for convenience, we use a Bayesian version due to Zhang (2006), although the “Generalized Fano” approach of Han and Verdú (1994) would also work. The bound in the high temperature regime is a “textbook” application that uses a uniform quadratic bound on the Kullback–Leibler (KL) divergence between the distributions determined by nearby parameter vectors. However, the moderate temperature regime uses a more refined analysis that does not appear to be standard. To facilitate the required computation, we use the Bregman divergence form of the KL divergence between Bernoulli distributions.
For the low temperature regime, this version of Fano’s inequality cannot be used, as the aforementioned KL divergence becomes unbounded. The basic form of Fano’s inequality is applicable, as the conditional entropy of the (randomly chosen) parameter given the data can be estimated. However, we were not able to obtain the optimal lower bound this way in this low temperature regime. Instead, we replicate the combinatorial argument of Long (1995) with modifications to handle finite $\beta$ (Theorem 2).
To establish upper bounds on the sample complexity, it is natural to consider the maximum likelihood estimator (MLE), which is equivalent to finding $\theta\in S^{d-1}$ that minimizes the empirical risk with respect to the logistic loss: $(1/n)\sum_{i=1}^{n}[g(\beta\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta)-\mathds{1}\{\mathbf{y}_{i}=1\}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta]$. The convexity of the logistic loss makes the MLE computationally tractable (perhaps after extending the parameter space to the ball). The risk of a given $\theta$ is the expected value of this empirical risk, and the excess risk relative to that of $\theta^{\star}$ is the KL divergence between the distributions determined by $\theta$ and $\theta^{\star}$. This KL divergence can, in turn, be related to the parameter error $\lVert\theta-\theta^{\star}\rVert$ using our analysis from the lower bound. However, bounding the excess risk sharply enough appears to be challenging. A standard approach in statistical learning theory is to use techniques like Rademacher averages from the theory of empirical processes to relate excess risks to the excess empirical risks. Unfortunately, using such tools designed specifically for smooth loss functions like the logistic loss (e.g., Srebro et al., 2010) leads to a sample complexity with suboptimal dependences on $\epsilon$ and $\beta$.
A more promising approach is to directly analyze the MLE by computing tight Taylor expansions of the estimation error in a way that keeps track of the dimension dependence (e.g., He and Shao, 2000; Portnoy, 1988), and, ideally, the inverse temperature. The most relevant work along these lines is that of Ostrovskii and Bach (2021) which, building on the fixed-design analysis of Bach (2010), leverages self-concordance properties of the logistic loss to directly bound the excess risk. [^2] The analysis of Bach (2010) and Ostrovskii and Bach (2021) can also apply to the setting where the logistic regression model is misspecified, but we only consider their analysis in the case where the model is well-specified. Although their analysis of the MLE matches the performance guarantees in the leading order terms as predicted by classical asymptotic analysis, it appears to be unable to avoid suboptimal dependence on the $\beta$ in the lower order terms.
We ultimately resort to establishing the upper bounds (Theorem 3), in all regimes, using empirical risk minimization with the classification loss, together with a precise analysis of the relationship between excess error rates and parameter distance. While this does not suggest any computationally tractable estimator (except, as far as we know, in the $\beta=\infty$ regime), the analysis is tight enough to match the lower bounds up to logarithmic factors in $1/\epsilon$.
Bounds on various Gaussian integrals used in our proofs are collected in Appendix A.
1.3 Other related works
Improper learning.
Several prior works analyze “improper learning” algorithms for (possibly misspecified) logistic regression (e.g., Kakade and Ng, 2004; Zhang, 2006; Hazan et al., 2014; Foster et al., 2018; Mourtada and Gaïffas, 2022) that do not necessarily produce an estimate of $\theta^{\star}$, which may not be sensible anyway if the model is misspecified. Instead, the goal of these algorithms is to achieve low prediction error guarantees.
Probit model misspecified as logistic.
As we were preparing this manuscript, we became aware of the recent work of Kuchelmeister and van de Geer (2023), which studies the estimation of the parameter vector $\theta^{\star}\in S^{d-1}$ in a probit regression model
$\mathbf{y}=\operatorname{sign}(\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}+\sigma\mathbf{z}),\quad(\mathbf{x},\mathbf{z})\sim\operatorname{N}(0,I_{d})\otimes\operatorname{N}(0,1),$
by maximizing the likelihood under a (misspecified) logistic regression model. Their analysis reveals two distinct regimes for the rate of convergence, depending on how the noise level $\sigma>0$ compares to the sample size and dimension. (They also consider estimation of the norm of the expected likelihood maximizer, which is related to the signal-to-noise ratio.) Although their data model and the high-level findings are related to those of our work, we note that the aim of our work is to understand (upper and lower bounds on) sample complexity in a well-specified setup, whereas Kuchelmeister and van de Geer study a specific estimator based on a misspecified model.
1.4 Notations
For notational convenience, we assume that a Bernoulli distribution $\operatorname{Bern}(\mu)$ has support on $\{-1,1\}$, with the mean parameter $\mu$ being the probability of $1$. We will, on occasion, associate each $\theta\in S^{d-1}$ with a homogeneous linear classifier $h_{\theta}\colon\mathbb{R}^{d}\to\{-1,1\}$, given by $h_{\theta}(x)=1$ if $x^{\scriptscriptstyle{\mathsf{T}}}\theta>0$ and $h_{\theta}(x)=-1$ otherwise. For any $\theta,\theta^{\prime}\in S^{d-1}$, let $\operatorname{err}_{\theta}(\theta^{\prime})=\Pr(h_{\theta^{\prime}}(\mathbf{x})\neq\mathbf{y})=\Pr(\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\prime}\leq 0)$ be the error rate of $h_{\theta^{\prime}}$ when the distribution of $(\mathbf{x},\mathbf{y})$ is specified by parameter $\theta$.
2 Lower bounds on the sample complexity
In this section, we give two lower bounds on the sample complexity.
2.1 Moderate and high temperatures
The following theorem establishes our sample complexity lower bound for moderate and high temperatures.
Theorem 1.
Fix $\epsilon\in\lparen 0,1\rparen$ ). Suppose ${\hat{\theta}}$ is an estimator that, for any $\theta^{\star}\in S^{d-1}$,
$\operatorname{\mathbb{E}}\left[\lVert{\hat{\theta}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})-\theta^{\star}\rVert\right]\leq\epsilon,$
where the expectation is with respect to the data with distribution determined by $\theta^{\star}$. Then the sample size $n$ must satisfy
$n\geq\frac{(d-1)\ln 2-\ln 4}{32\epsilon^{2}\beta\min\{\beta,2\sqrt{2/\pi}\}}.$
The proof of Theorem 1 uses the following information theoretic lower method due to Zhang (2006).
Lemma 1 (Theorem 6.1 of Zhang, 2006).
Let $\Pi$ be a probabiltiy measure on a parameter space $\Theta$ indexing a family of probability measures $(P_{\theta})_{\theta\in\Theta}$ on a data space $\mathcal{Z}$, and let $L\colon\Theta\times\Theta\to\mathbb{R}$ be a loss function. Let $\boldsymbol{\theta}\sim\Pi$ and $\mathbf{Z}\mid\boldsymbol{\theta}\sim P_{\boldsymbol{\theta}}$. For any (possibly randomized) estimator ${\hat{\theta}}\colon\mathcal{Z}\to\Theta$,
$\operatorname{\mathbb{E}}\left[L(\boldsymbol{\theta},{\hat{\theta}}(\mathbf{Z}))\right]\geq\frac{1}{2}\sup\left\{\varepsilon:\inf_{\theta\in\Theta}-\ln(\Pi(B(\theta,\varepsilon)))\geq 2\kappa+\ln 4\right\}$
where $B(\theta,\varepsilon)=\{\theta^{\prime}\in\Theta:L(\theta,\theta^{\prime})<\varepsilon\}$ and $\kappa=\operatorname{\mathbb{E}}_{(\boldsymbol{\theta},\boldsymbol{\theta}^{\prime})\sim\Pi\otimes\Pi}\left[\operatorname{KL}(P_{\boldsymbol{\theta}}\|P_{\boldsymbol{\theta}^{\prime}})\right]$.
Proof of Theorem 1.
We prove the contrapositive. Assume that
$n<\frac{(d-1)\ln 2-\ln 4}{32\epsilon^{2}\beta\min\{\beta,2\sqrt{2/\pi}\}}.$
Set $\delta=4\epsilon$. Fix a unit vector $u\in S^{d-1}$, and let $C=\{\theta\in S^{d-1}:\lVert\theta-u\rVert\leq\delta\}$ be the spherical cap of radius $\delta$ around $u$. Let $\Theta=S^{d-1}$, $\Pi$ be the uniform measure on $C$, and $L(\theta,\theta^{\prime})=\lVert\theta-\theta^{\prime}\rVert$. For each $\theta\in\Theta$, we let $P_{\theta}$ denote the joint distribution of the data $\mathbf{Z}=((\mathbf{x}_{1},\mathbf{y}_{1}),\dotsc,(\mathbf{x}_{n},\mathbf{y}_{n}))$ as determined by our model with parameter $\theta$. Observe that $\Pi(B(\theta,\varepsilon))\leq(\varepsilon/\delta)^{d-1}$ for any $\varepsilon\leq\delta$. Therefore, provided that
$(d-1)\ln 2>2\kappa+\ln 4,$
we have $\operatorname{\mathbb{E}}[\lVert\boldsymbol{\theta}-{\hat{\theta}}(\mathbf{Z})\rVert]>\delta/4=\epsilon$ by Lemma 1. So it remains to establish the above displayed inequality.
Since the $n$ data are i.i.d. in each of $P_{\theta}$ and $P_{\theta^{\prime}}$, and the marginal distribution of $\mathbf{x}$ is the same in both $P_{\theta}$ and $P_{\theta^{\prime}}$, the chain rule for KL divergence implies
$\operatorname{KL}(P_{\theta}\|P_{\theta^{\prime}})=n\operatorname{\mathbb{E}}\left[\operatorname{KL}(\operatorname{Bern}(g^{\prime}(\beta\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta))\|\operatorname{Bern}(g^{\prime}(\beta\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\prime})))\right].$
Let $\mathbf{z}=\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta$ and $\mathbf{z}^{\prime}=\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\prime}$, so each of $\mathbf{z}$ and $\mathbf{z}^{\prime}$ is a standard normal random variable, and the correlation $\rho$ between $\mathbf{z}$ and $\mathbf{z}^{\prime}$ satisfies
$\rho=\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\prime}=1-\frac{1}{2}\lVert\theta-\theta^{\prime}\rVert^{2}\geq 1-\frac{1}{2}(2\delta)^{2}=1-2\delta^{2},$
where we have used the triangle inequality $\lVert\theta-\theta^{\prime}\rVert\leq\lVert\theta-u\rVert+\lVert\theta^{\prime}-u\rVert\leq 2\delta$. By Lemma 10,
$\operatorname{\mathbb{E}}\left[\operatorname{KL}(\operatorname{Bern}(g^{\prime}(\beta\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta))\|\operatorname{Bern}(g^{\prime}(\beta\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\prime})))\right]$ $=\operatorname{\mathbb{E}}\left[\operatorname{KL}(\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}))\|\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}^{\prime})))\right]$
$\leq\frac{\beta}{2}\lparen 1-\rho\rparen\min\left\{\beta,2\sqrt{2/\pi}\right\}$
$\leq\delta^{2}\beta\min\left\{\beta,2\sqrt{2/\pi}\right\}.$
Therefore
$2\kappa+\ln 4\leq 2n\delta^{2}\beta\min\{\beta,2\sqrt{2/\pi}\}+\ln 4<(d-1)\ln 2$
as required, where the second inequality uses the assumption on $n$. ∎
2.2 Low temperatures
The sample complexity lower bound from Theorem 1 tends to 00 as $\beta\to\infty$. This appears to be a limitation of the proof technique, which is only useful for $\beta\lesssim 1/\epsilon$. We next establish a lower bound that improves on Theorem 1 for $\beta\gg 1/\epsilon$.
Theorem 2.
Fix $\epsilon\in\lparen 0,1\rparen$ ). Assume $\beta\geq 4\sqrt{2/\pi}/\epsilon$. Suppose ${\hat{\theta}}$ is an estimator that, for any $\theta^{\star}\in S^{d-1}$,
$\Pr\lparen\lVert{\hat{\theta}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})-\theta^{\star}\rVert<\epsilon\rparen\geq\frac{1}{2},$
where the probability is with respect to the data with distribution determined by $\theta^{\star}$. Then the sample size $n$ must satisfy
$n\geq\frac{d-1}{\epsilon}\cdot\frac{\log(\log(2e/\epsilon))-\log(16e)}{(1+o(1))\log(2e/\epsilon)}$
where the $o(1)$ term depends only $\epsilon$ and vanishes as $\epsilon\to 0$. If $\beta=\infty$, then the sample size must, in fact, satisfy $n\geq(d-1)/(8e\epsilon)$.
The proof of Theorem 2 essentially follows that of Long (1995) for a lower bound on the sample complexity of PAC learning homogeneous linear classifiers under the uniform distribution on $S^{d-1}$. In Long’s setting, the data distribution coincides with ours for $\beta=\infty$. We make a minor modification to the argument to also handle finite (but large) $\beta$. The extra $\log\log(1/\epsilon)/\log(1/\epsilon)$ factor in the finite $\beta$ case appears to be an artifact of the proof technique. We also note that the $\beta=\infty$ case is directly implied by the main result of Long (1995).
Long’s proof, as well as ours, relies on the following bound on the shattering number of homogeneous linear classifiers.
Lemma 2 (Winder, 1966, Corollary on page 816).
Let $H$ be the family of homogeneous linear classifiers in $\mathbb{R}^{d}$. For any $x_{1},\dotsc,x_{n}\in\mathbb{R}^{d}$,
|$\lvert\{(h(x_{1}),\dotsc,h(x_{n})):h\in H\}\rvert\leq 2\left\lparen\frac{ne}{d-1}\right\rparen^{d-1}.$
We also need the following well-known lower bound on the packing number for $S^{d-1}$; the proof is given for completeness.
Lemma 3.
There exists an $\epsilon$-packing of the unit sphere $S^{d-1}$ with respect to $\ell_{2}$ distance of cardinality $(1/\epsilon)^{d-1}$.
Proof.
Let $W$ be a subspace of $d$-dimensional Euclidean space of dimension $d-1$. By a standard volume argument, there is an $\epsilon$-packing $p_{1},\dotsc,p_{M}$ of the unit ball in $W$ with $M\geq(1/\epsilon)^{d-1}$. For each point $p_{i}$, there is a corresponding point $p_{i}^{\prime}\in S^{d-1}$ whose orthogonal projection to $W$ is $p_{i}$. For any $i\neq j$, we have $\lVert p_{i}^{\prime}-p_{j}^{\prime}\rVert\geq\lVert p_{i}-p_{j}\rVert\geq\epsilon$, so $p_{1}^{\prime},\dotsc,p_{M}^{\prime}$ is an $\epsilon$-packing of $S^{d-1}$. ∎
Finally, we need the following bound on the error rate of the homogeneous linear classifier $h_{\theta^{\star}}$. (The lemma holds for all $\beta\geq 0$, not just $\beta\gtrsim 1/\epsilon$.)
Lemma 4.
The error rate of $h_{\theta^{\star}}$ satisfies
$\operatorname{err}_{\theta^{\star}}(\theta^{\star})\leq\min\left\{\frac{1}{2},\ \frac{1}{2}-\frac{\beta}{4}\sqrt{\frac{2}{\pi}}\left\lparen 1-\frac{\beta^{2}}{6}\right\rparen,\ \frac{1}{\beta}\sqrt{\frac{2}{\pi}}\right\}.$
Proof.
Since
$\operatorname{err}_{\theta^{\star}}(\theta^{\star})=\Pr(\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0)=\operatorname{\mathbb{E}}[g^{\prime}(-\beta\lvert\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\rvert)]=\operatorname{\mathbb{E}}[g^{\prime}(-\beta\lvert\mathbf{z}\rvert)]$
for a normal random variable $\mathbf{z}$, the claim follows immediately from Lemma 9. ∎
Proof of Theorem 2.
Assume without loss of generality that $\epsilon\leq 1/2$. Let $\mathbf{x}_{1},\dotsc,\mathbf{x}_{n}$ be independent standard normal random vectors. Let $\mathbf{y}_{1}^{\theta},\dotsc,\mathbf{y}_{n}^{\theta}$ for all $\theta\in S^{d-1}$ be independent Bernoulli random variables with $\mathbf{y}_{i}^{\theta}\mid(\mathbf{x}_{1},\dotsc,\mathbf{x}_{n})\sim\operatorname{Bern}(g^{\prime}(\beta\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta))$, and let $\mathbf{Z}^{\theta}=((\mathbf{x}_{1},\mathbf{y}_{1}^{\theta}),\dotsc,(\mathbf{x}_{n},\mathbf{y}_{n}^{\theta}))$. So the data $\mathbf{Z}^{\theta}$ follows our model with parameter $\theta$, but all $\{\mathbf{Z}^{\theta}:\theta\in S^{d-1}\}$ share the same $\mathbf{x}_{1},\dotsc,\mathbf{x}_{n}$. Let $\operatorname{\widehat{err}}_{\theta}(\theta)=(1/n)\sum_{i=1}^{n}\mathds{1}\{h_{\theta}(\mathbf{x}_{i})\neq\mathbf{y}_{i}^{\theta}\}$ be the empirical error rate of $h_{\theta}$ (with respect to $\mathbf{Z}^{\theta}$).
Suppose $\Pr(\lVert{\hat{\theta}}(\mathbf{Z}^{\theta})-\theta\rVert<\epsilon)\geq 1/2$ for all $\theta\in S^{d-1}$. By Lemma 4, since $\beta\geq 4\sqrt{2/\pi}/\epsilon$, we have
$\operatorname{err}_{\theta}(\theta)\leq\frac{\epsilon}{4}.$
Let $U$ be a $2\epsilon$-packing of $S^{d-1}$ with respect to $\ell_{2}$ of cardinality $\lvert U\rvert\geq(2\epsilon)^{-(d-1)}$, as guaranteed to exist by Lemma 3. Let $G_{\theta}$ be the (indicator of) the event $\lVert{\hat{\theta}}(\mathbf{Z}^{\theta})-\theta\rVert<\epsilon$ and $\operatorname{\widehat{err}}_{\theta}(\theta)\leq\epsilon$. By Markov’s inequality,
$\Pr(\operatorname{\widehat{err}}_{\theta}(\theta)\geq\epsilon)\leq\frac{1}{4}.$
So
$\operatorname{\mathbb{E}}\left[\sum_{\theta\in U}G_{\theta}\right]\geq\sum_{\theta\in U}\operatorname{\mathbb{E}}[G_{\theta}]\geq\sum_{\theta\in U}\left\lparen 1-\frac{1}{2}-\frac{1}{4}\right\rparen=\frac{\lvert U\rvert}{4}\geq\frac{1}{4}\left\lparen\frac{1}{2\epsilon}\right\rparen^{d-1}.$
Consider any two $\theta,\theta^{\prime}\in U$. If $\mathbf{Z}^{\theta}=\mathbf{Z}^{\theta^{\prime}}$, then either $\lVert{\hat{\theta}}(\mathbf{Z}^{\theta})-\theta\rVert>\epsilon$ or $\lVert{\hat{\theta}}(\mathbf{Z}^{\theta^{\prime}})-\theta^{\prime}\rVert>\epsilon$ (since $\lVert\theta-\theta^{\prime}\rVert\geq 2\epsilon$), so at least one of $G_{\theta}$ and $G_{\theta^{\prime}}$ is zero. Moreover, if $G_{\theta}=1$, then the labels $\mathbf{y}_{i}^{\theta}$ are realized by the homogeneous linear classifier determined by $\theta$—except for up to $\lfloor n\epsilon\rfloor$ labels, which could be flipped. Therefore $\operatorname{\mathbb{E}}\left[\sum_{\theta\in U}G_{\theta}\right]$ is at most the number of ways of labeling $(\mathbf{x}_{1},\dotsc,\mathbf{x}_{n})$ by homogeneous linear classifiers determined by weight vectors from $U$, multiplied by $2^{\lfloor n\epsilon\rfloor}\binom{n}{\lfloor n\epsilon\rfloor}$. Hence, by Lemma 2,
[…]
where $M=(d-1)/(n\epsilon)$. Combining the upper and lower bounds on $\operatorname{\mathbb{E}}\left[\sum_{\theta\in U}G_{\theta}\right]$ gives
$\frac{e}{M\epsilon}\cdot\left\lparen\frac{2e}{\epsilon}\right\rparen^{1/M}\geq\frac{1}{8^{1/(d-1)}}\cdot\frac{1}{2\epsilon}\geq\frac{1}{16\epsilon}.$
Taking logarithms of both sides and simplifying gives the following inequality:
[…]
Let $T(\epsilon)=\log(2e/\epsilon)/(16e)$. Then, using asymptotic expansion of the product log function (Corless et al., 1996), we have
$\frac{M}{16e}\leq\frac{T(\epsilon)}{\log T(\epsilon)}(1+o(1)),$
where the $o(1)$ term vanishes as $T(\epsilon)\to\infty$ (i.e., $\epsilon\to 0$). This implies
$n\geq\frac{d-1}{16e\epsilon}\cdot\frac{\log T(\epsilon)}{(1+o(1))T(\epsilon)}$
as claimed.
If $\beta=\infty$, then $\Pr(\operatorname{\widehat{err}}_{\theta}(\theta)=0)=1$, and hence we have
[…]
Therefore
$n\geq\frac{d-1}{8e\epsilon}.\qed$
3 Upper bounds on the sample complexity
We complement the sample complexity lower bounds from Theorems 1 and 2 with the following upper bound.
Theorem 3.
There is an absolute constant $C>0$ and an estimator ${\hat{\theta}}$ such that, for any $\theta^{\star}\in S^{d-1}$ and any $\epsilon,\delta\in\lparen 0,1\rparen$ ), we have
$\Pr\lparen\lVert{\hat{\theta}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})-\theta^{\star}\rVert\leq\epsilon\rparen\geq 1-\delta,$
where the probability is with respect to the data with distribution determined by $\theta^{\star}$, provided that
$n\geq C\min\left\{\left\lparen\frac{d\log(1/\varepsilon)+\log(1/\delta)}{\beta\varepsilon^{2}}+\frac{d\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon}\right\rparen,\frac{d+\log(1/\delta)}{\beta^{2}\varepsilon^{2}}\right\}.$
The estimator referenced in Theorem 3 is based on empirical risk minimization. We describe this estimator and its properties below. Fix a “ground truth” parameter $\theta^{\star}\in S^{d-1}$ determining the distribution of $(\mathbf{x},\mathbf{y})$, and note that
$\min_{\theta\in S^{d-1}}\operatorname{err}_{\theta^{\star}}(\theta)=\operatorname{err}_{\theta^{\star}}(\theta^{\star}).$
The empirical risk minimizer ${\hat{\theta}_{\operatorname{ERM}}}={\hat{\theta}_{\operatorname{ERM}}}((\mathbf{x}_{i},\mathbf{y}_{i})_{i=1}^{n})$ is the $\theta\in S^{d-1}$ that minimizes
$\frac{1}{n}\sum_{i=1}^{n}\mathds{1}\{\mathbf{y}_{i}\mathbf{x}_{i}^{\scriptscriptstyle{\mathsf{T}}}\theta\leq 0\}.$
(If the minimizer not unique, we choose one arbitrarily.) The following is a standard performance guarantee for empirical risk minimization, combined with the fact that the VC dimension of the class of homogeneous linear classifiers is $d$ [^3] Note that Lemma 5 is true under every distribution for $(\mathbf{x},\mathbf{y})$; it does not rely on specific properties of our model. In the special case of where $\mathbf{x}$ is the $d$-dimensional standard normal (or any other spherically symmetric distribution) and $\beta=\infty$ (which implies $\operatorname{opt}=0$), the $\log(1/\varepsilon)$ can be removed (Long, 1995). So, in this case, the sample size requirement is $C(d+\log(1/\delta))/\varepsilon$.
Lemma 5 (Vapnik and Chervonenkis, 1971; Talagrand, 1994).
There is a universal constant $C>0$ such that the following holds. Let $\operatorname{opt}=\min_{\theta\in S^{d-1}}\operatorname{err}_{\theta^{\star}}(\theta)$. For any $\varepsilon\in\lparen 0,1\rparen$ ) and $\delta\in\lparen 0,1\rparen$ ), if
$n\geq C\min\left\{\left\lparen\frac{d\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon^{2}}\operatorname{opt}+\frac{d\log(1/\varepsilon)+\log(1/\delta)}{\varepsilon}\right\rparen,\frac{d+\log(1/\delta)}{\varepsilon^{2}}\right\},$
then the empirical risk minimizer ${\hat{\theta}_{\operatorname{ERM}}}$ satisfies
$\operatorname{err}_{\theta^{\star}}({\hat{\theta}_{\operatorname{ERM}}})-\operatorname{opt}\leq\varepsilon$
with probability at least $1-\delta$ over the realization of the data.
The following lemma is used to relate $\lVert{\hat{\theta}_{\operatorname{ERM}}}-\theta^{\star}\rVert$ to $\operatorname{err}_{\theta^{\star}}({\hat{\theta}_{\operatorname{ERM}}})-\operatorname{err}_{\theta^{\star}}(\theta^{\star})$.
Lemma 6.
If $\theta\in S^{d-1}$ satisfies
$\operatorname{err}_{\theta^{\star}}(\theta)-\operatorname{err}_{\theta^{\star}}(\theta^{\star})\leq\varepsilon,$
then
$\lVert\theta-\theta^{\star}\rVert\leq\frac{\pi\varepsilon}{1-2\operatorname{err}_{\theta^{\star}}(\theta^{\star})}.$
Proof.
The distribution of $\mathbf{y}$ depends on $\mathbf{x}$ only through $\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}$. Therefore, since $\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta$ and $\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}$ are each standard normal random variables with correlation $\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}$, Lemma 11 implies
$\Pr(\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0\mid\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}>0)=\frac{\arccos(\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})}{\pi}=1-\Pr(\theta^{\scriptscriptstyle{\mathsf{T}}}\mathbf{x}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}>0\mid\mathbf{y}\mathbf{x}^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star}\leq 0).$
Therefore, the error rate of $\theta$ is
$\operatorname{err}_{\theta^{\star}}(\theta)=\frac{\arccos(\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})}{\pi}\left\lparen 1-\operatorname{err}_{\theta^{\star}}(\theta^{\star})\right\rparen+\left\lparen 1-\frac{\arccos(\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})}{\pi}\right\rparen\operatorname{err}_{\theta^{\star}}(\theta^{\star}),$
and hence
$\operatorname{err}_{\theta^{\star}}(\theta)-\operatorname{err}_{\theta^{\star}}(\theta^{\star})=\frac{\arccos(\theta^{\scriptscriptstyle{\mathsf{T}}}\theta^{\star})}{\pi}\left\lparen 1-2\operatorname{err}_{\theta^{\star}}(\theta^{\star})\right\rparen\geq\frac{\lVert\theta-\theta^{\star}\rVert}{\pi}\left\lparen 1-2\operatorname{err}_{\theta^{\star}}(\theta^{\star})\right\rparen$ ) )
where the inequality uses $\arccos(\rho)\geq\sqrt{2(1-\rho)}$. Combining this inequality with the hypothesis that $\operatorname{err}_{\theta^{\star}}(\theta)-\operatorname{err}_{\theta^{\star}}(\theta^{\star})\leq\varepsilon$ completes the proof. ∎
Finally, to deduce upper bounds on the sample complexity for achieving $\lVert{\hat{\theta}_{\operatorname{ERM}}}-\theta^{\star}\rVert\leq\epsilon$, we combine Lemma 5 and Lemma 6 with the bound on $\operatorname{opt}=\operatorname{err}_{\theta^{\star}}(\theta^{\star})$ from Lemma 4.
4 Discussion
Our characterization of the sample complexity of estimation in logistic regression deliniates the high, moderate, and low temperature regimes. However, we are only aware of computationally tractable estimators that achieve the (near) optimal sample complexities at high and zero temperatures: e.g., the linear estimator of Plan and Vershynin (2012) for $\beta\lesssim 1$, and the estimator based on solving a linear feasibility program (or the algorithm of Balcan and Long (2013)) for $\beta=\infty$. It would be interesting to determine if MLE achieves optimal sample complexity in the other regimes, and if the method of optimization is important—for instance, due to the implicit biases of the optimization procedure (Telgarsky, 2013; Soudry et al., 2018; Ji and Telgarsky, 2019).
Appendix A Bounds on Gaussian integrals
Lemma 7 (Feller, 1968, page 175).
Let $\mathbf{z}\sim\operatorname{N}(0,1)$. For any $t>0$,
$\Pr(\mathbf{z}\geq t)\leq\frac{1}{t\sqrt{2\pi}}\exp(-t^{2}/2).$
Recall that $g(\eta)=\ln(1+\exp(\eta))$.
Lemma 8.
Let $\mathbf{z}\sim\operatorname{N}(0,1)$. For any $\beta>0$,
$\operatorname{\mathbb{E}}[g^{\prime\prime}(\beta\mathbf{z})]\leq\min\left\{\frac{1}{2},\ \operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen]\right\}$
and
$\operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen]\leq\frac{1}{\beta}\sqrt{\frac{2}{\pi}}.$
Proof.
First, observe that
$g^{\prime\prime}(\eta)=\frac{1}{(1+\exp(\eta))(1+\exp(-\eta))}\leq\min\left\{\frac{1}{2},\ \exp(-\lvert\eta\rvert)\right\}.$
This establishes the first inequality. For the second inequality,
$\operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen]$ $=2\int_{0}^{\infty}\frac{1}{\sqrt{2\pi}}\exp\lparen-\beta z-z^{2}/2\rparen\operatorname{d\!}z$
$=2\exp\lparen\beta^{2}/2\rparen\int_{0}^{\infty}\frac{1}{\sqrt{2\pi}}\exp\lparen-(z+\beta)^{2}/2\rparen\operatorname{d\!}z$
$=2\exp\lparen\beta^{2}/2\rparen\Pr(\mathbf{z}\geq\beta)$ )
$\leq\frac{1}{\beta}\sqrt{\frac{2}{\pi}},$
where the inequality follows by Lemma 7. ∎
Remark 1.
Using a similar analysis as above, one can check that the bounds in Lemma 8 are fairly tight up to constants:
$\operatorname{\mathbb{E}}[g^{\prime\prime}(\beta\mathbf{z})]\geq\frac{1}{4}\operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen]$
and
[…]
Lemma 9.
Let $\mathbf{z}\sim\operatorname{N}(0,1)$. For any $\beta>0$,
$\operatorname{\mathbb{E}}[g^{\prime}(-\beta\lvert\mathbf{z}\rvert)]\leq\min\left\{\frac{1}{2},\ \frac{1}{2}-\frac{\beta}{4}\sqrt{\frac{2}{\pi}}\left\lparen 1-\frac{\beta^{2}}{6}\right\rparen,\ \frac{1}{\beta}\sqrt{\frac{2}{\pi}}\right\}.$
Proof.
First, observe that $g^{\prime}(\eta)\leq 1/2$ for all $\eta\leq 0$. Furthermore, by Taylor’s theorem, for any $\eta\in\mathbb{R}$, there exists $h\in\mathbb{R}$ (between 00 and $\eta$) such that
$g^{\prime}(\eta)$ $=g^{\prime}(0)+g^{\prime\prime}(\eta)\eta+\frac{1}{2}g^{\prime\prime\prime}(\eta)\eta^{2}+\frac{1}{6}g^{\prime\prime\prime\prime}(h)\eta^{3}$
$\leq g^{\prime}(0)+g^{\prime\prime}(0)\eta+\frac{1}{2}g^{\prime\prime\prime}(0)\eta^{2}+\frac{1}{48}\lvert\eta\rvert^{3}$
$=\frac{1}{2}+\frac{1}{4}\eta+\frac{1}{48}\lvert\eta\rvert^{3},$
where the inequality uses the fact that $\lvert g^{\prime\prime\prime\prime}(h)\rvert\leq 1/8$ for all $h\in\mathbb{R}$. Therefore
$\operatorname{\mathbb{E}}[g^{\prime}(-\beta\lvert\mathbf{z}\rvert)]$ $\leq\frac{1}{2}-\frac{\beta}{4}\operatorname{\mathbb{E}}[\lvert\mathbf{z}\rvert]+\frac{\beta^{3}}{48}\operatorname{\mathbb{E}}[\lvert\mathbf{z}\rvert^{3}]$
$=\frac{1}{2}-\frac{\beta}{4}\sqrt{\frac{2}{\pi}}+\frac{\beta^{3}}{24}\sqrt{\frac{2}{\pi}}$
$=\frac{1}{2}-\frac{\beta}{4}\sqrt{\frac{2}{\pi}}\left\lparen 1-\frac{\beta^{2}}{6}\right\rparen.$ ) .
Finally, we also have
$\operatorname{\mathbb{E}}[g^{\prime}(-\beta\lvert\mathbf{z}\rvert)]=\operatorname{\mathbb{E}}\left[\frac{1}{1+\exp\lparen\beta\lvert\mathbf{z}\rvert\rparen}\right]\leq\operatorname{\mathbb{E}}[\exp\lparen-\beta\lvert\mathbf{z}\rvert\rparen]\leq\frac{1}{\beta}\sqrt{\frac{2}{\pi}},$
where the final inequality follows by Lemma 8. ∎
Recall that the Bernoulli distributions form an exponential family $\{p_{\eta}:\eta\in\mathbb{R}\}$, where
$p_{\eta}(y)=\exp\left\lparen\eta\mathds{1}\{y=1\}-g(\eta)\right\rparen$ )
and $g(\eta)=\ln(1+\exp(\eta))$ is the log partition function for $p_{\eta}$. The mean parameter—i.e., the mean of $\mathds{1}\{y=1\}$ under $p_{\eta}$—is given by $g^{\prime}(\eta)$. In the proof of the following lemma, we use the fact that the KL divergence $\operatorname{KL}(p_{\eta}\|p_{\eta^{\prime}})$ can be expressed as a Bregman divergence associated with $g$:
$\operatorname{KL}(p_{\eta}\|p_{\eta^{\prime}})=g(\eta^{\prime})-\left\lparen g(\eta)+g^{\prime}(\eta)(\eta^{\prime}-\eta)\right\rparen.$ ) .
Lemma 10.
Let $\mathbf{z}$ and $\mathbf{z}^{\prime}$ be $\operatorname{N}(0,1)$ random variables with correlation $\rho$. Then
$\operatorname{\mathbb{E}}\left[\operatorname{KL}(\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}))\|\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}^{\prime})))\right]\leq\frac{\beta}{2}(1-\rho)\min\left\{\beta,2\sqrt{2/\pi}\right\}.$
Proof.
For any $\eta,\eta^{\prime}\in\mathbb{R}$, we have
$\operatorname{KL}(\operatorname{Bern}(g^{\prime}(\eta))\|\operatorname{Bern}(g^{\prime}(\eta^{\prime})))=g(\eta^{\prime})-g(\eta)-g^{\prime}(\eta)\lparen\eta^{\prime}-\eta\rparen.$ ) .
Hence,
$\operatorname{\mathbb{E}}[\operatorname{KL}(\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}))\|\operatorname{Bern}(g^{\prime}(\beta\mathbf{z}^{\prime})))]$ $=\operatorname{\mathbb{E}}[g(\beta\mathbf{z}^{\prime})-g(\beta\mathbf{z})-g^{\prime}(\beta\mathbf{z})\beta(\mathbf{z}^{\prime}-\mathbf{z})]$
$=\operatorname{\mathbb{E}}[g^{\prime}(\beta\mathbf{z})\beta\lparen\mathbf{z}-\mathbf{z}^{\prime}\rparen].$
Since $\mathbf{z}$ and $\mathbf{z}^{\prime}$ have correlation $\rho$, we may write
$\mathbf{z}^{\prime}=\rho\mathbf{z}+\sqrt{1-\rho^{2}}\mathbf{z}_{\perp},$
where $\mathbf{z}_{\perp}\sim\operatorname{N}(0,1)$ is independent of $\mathbf{z}$. Thus
$\operatorname{\mathbb{E}}[g^{\prime}(\beta\mathbf{z})\beta\lparen\mathbf{z}-\mathbf{z}^{\prime}\rparen]$ $=\operatorname{\mathbb{E}}\left[g^{\prime}(\beta\mathbf{z})\beta\lparen 1-\rho\rparen\mathbf{z}-\sqrt{1-\rho^{2}}\mathbf{z}_{\perp}\right]$
$=\beta\lparen 1-\rho\rparen\operatorname{\mathbb{E}}[g^{\prime}(\beta\mathbf{z})\mathbf{z}]$
$=\beta^{2}\lparen 1-\rho\rparen\operatorname{\mathbb{E}}[g^{\prime\prime}(\beta\mathbf{z})]$
$\leq\frac{\beta}{2}\lparen 1-\rho\rparen\min\left\{\beta,2\sqrt{2/\pi}\right\},$
where the penultimate step follows by Stein’s identity, and the final step follows by Lemma 8. ∎
Lemma 11.
Let $\mathbf{z}$ and $\mathbf{z}^{\prime}$ be $\operatorname{N}(0,1)$ random variables with correlation $\rho$. Then
$\Pr(\mathbf{z}\mathbf{z}^{\prime}\leq 0\mid\mathbf{z}<0)=\Pr(\mathbf{z}\mathbf{z}^{\prime}\leq 0\mid\mathbf{z}>0)=\frac{\arccos(\rho)}{\pi}.$
Proof.
The claim is trivially true if $\lvert\rho\rvert=1$ since $\arccos(1)/\pi=0$ and $\arccos(-1)/\pi=1$. So without loss of generality assume $\lvert\rho\rvert<1$. Since $\mathbf{z}$ and $\mathbf{z}^{\prime}$ have correlation $\rho$, we may write
$\mathbf{z}^{\prime}=\rho\mathbf{z}+\sqrt{1-\rho^{2}}\mathbf{z}_{\perp},$
where $\mathbf{z}_{\perp}\sim\operatorname{N}(0,1)$ is independent of $\mathbf{z}$. Then, with $m=\rho/\sqrt{1-\rho^{2}}$,
$\Pr\lparen\mathbf{z}\mathbf{z}^{\prime}\leq 0\mid\mathbf{z}<0\rparen=\Pr\lparen\mathbf{z}_{\perp}>-m\mathbf{z}\mid\mathbf{z}<0\rparen.$ ) .
Because the joint distribution of $(\mathbf{z},\mathbf{z}_{\perp})$ is spherically symmetric, the probability in the previous display is equal to the proportion of the “left half” ($x<0$) of the unit circle in the $x$-$y$ plane that lies above the $y=-mx$ line: this proportion is $\arccos(\rho)/\pi$. An analogous argument proves the same identity for $\Pr(\mathbf{z}\mathbf{z}^{\prime}\leq 0\mid\mathbf{z}>0)$. ∎